OpenStudy (anonymous):

f(3)= -5, g(3)=9, f'(3)=8 and g'(3)= -4 how do I find (fg)'(3) (f/g)'(3) and (f/(f-g))'(3)??

OpenStudy (anonymous):

use the rules for derivatives of products and quotients for example, the simplest one is \[(fg)'=f'g+g'f=8\times 9+(-4)\times (-5)\] others are similar. rules plus arithmetic